OpenStudy (anonymous):

Given: Triangle ABC is similar to triangle DEF. AB = 8, BC = 12, EF = 9, DE = 6 What is the scale factor of triangle ABC to triangle DEF

OpenStudy (anonymous):

a) 12/9 b) 8/6 c) 4/3 d) all of the above

OpenStudy (anonymous):

12/9 = 8/6 = 4/3 Hence (regardless of the question itself)... d) all of the above

OpenStudy (anonymous):

d